# Sound Unblending: Exploring Sound Manipulations for Accessible Mixed-Reality Awareness

### v1

@@ Abstract.
Mixed-reality (MR) soundscapes blend real-world sound with virtual audio from hearing devices, presenting intricate auditory information that is hard to discern and differentiate. This is particularly challenging for blind or visually impaired individuals, who rely on sounds and descriptions in their everyday lives. To understand how complex audio information is consumed, we analyzed online forum posts within the blind community, identifying prevailing challenges, needs, and desired solutions. We synthesized the results and proposed Sound Unblending for increasing MR sound awareness, which includes six sound manipulations: Ambience Builder, Feature Shifter, Earcon Generator, Prioritizer, Spatializer, and Stylizer. To evaluate the effectiveness of sound unblending, we conducted a user study with 18 blind participants across three simulated MR scenarios, where participants identified specific sounds within intricate soundscapes. We found that sound unblending increased MR sound awareness and minimized cognitive load. Finally, we developed three real-world example applications to demonstrate the practicality of sound unblending.
@@
copyright: rightsretainedjournalyear: 2024conference: ; January; 2024price: 15.00isbn: XXX-X-XXXX-XXXX-X/XX/XXccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Virtual reality
@@
Figure 1. We present Sound Unblending, a concept to manipulate sounds for improving mixed-reality awareness. (a) Sound Unblending situates in the auditory Reality-Virtuality Continuum with full transparency and noise cancellation as two ends, and comprises (b) six sound manipulators, which are Ambience Builder, Feature Shifter, Stylizer, Spatializer, Earcon Generator, and Prioritizer. (c) In a scenario, a BVI user navigates a busy street with a white cane and audio directions. sometimes may receive ringtones and pass by construction sites with drilling noises. (d.1) Ambience Builder makes the auditory transparency half to suppress nuanced noises while retaining real-world awareness. (d.2) Feature Shifter increases the white cane sounds to make them distinctive. (d.3) Earcon Generator plays an earcon to signal the danger, and Stylizer applies a low-pass filter to make drilling noise less sharp to hear. (d.4) Prioritizer delays the audio directions when they conflict with drilling noises. (d.5) Spatializer places ringtone on the right and audio directions on the left to increase distinguishability.
We present Sound Unblending, a concept to manipulate sounds for improving mixed-reality awareness. (a) Sound Unblending situates in the auditory Reality-Virtuality Continuum with full transparency and noise cancellation as two ends, and comprises (b) six sound manipulators, which are Ambience Builder, Feature Shifter, Stylizer, Spatializer, Earcon Generator, and Prioritizer. (c) In a scenario, a BVI user navigates a busy street with a white cane and audio directions. sometimes may receive ringtones and pass by construction sites with drilling noises. (d.1) Ambience Builder makes the auditory transparency half to suppress nuanced noises while retaining real-world awareness. (d.2) Feature Shifter increases the white cane sounds to make them distinctive. (d.3) Earcon Generator plays an earcon to signal the danger, and Stylizer applies a low-pass filter to make drilling noise less sharp to hear. (d.4) Prioritizer delays the audio directions when they conflict with drilling noises. (d.5) Spatializer places ringtone on the right and audio directions on the left to increase distinguishability.
Figure 1. We present Sound Unblending, a concept to manipulate sounds for improving mixed-reality awareness. (a) Sound Unblending situates in the auditory Reality-Virtuality Continuum with full transparency and noise cancellation as two ends, and comprises (b) six sound manipulators, which are Ambience Builder, Feature Shifter, Stylizer, Spatializer, Earcon Generator, and Prioritizer. (c) In a scenario, a BVI user navigates a busy street with a white cane and audio directions. sometimes may receive ringtones and pass by construction sites with drilling noises. (d.1) Ambience Builder makes the auditory transparency half to suppress nuanced noises while retaining real-world awareness. (d.2) Feature Shifter increases the white cane sounds to make them distinctive. (d.3) Earcon Generator plays an earcon to signal the danger, and Stylizer applies a low-pass filter to make drilling noise less sharp to hear. (d.4) Prioritizer delays the audio directions when they conflict with drilling noises. (d.5) Spatializer places ringtone on the right and audio directions on the left to increase distinguishability.
@@ 1. Introduction
To understand the current practices in consuming complex audio information, we first conducted a content analysis from active online forums within the BVI community, where we collected posts and comments about the prevailing scenarios, challenges, and potential solutions for consuming complex audio information. We found that many scenarios require the consumption of RW [^1] RW sounds refer to sounds stemming from RW environments, such as crowd, speaker, television. and VR [^2] VR sounds refer to sounds rendered through users’ hearing devices. sounds, such as navigating a busy street with audio directions, playing an instrument following music tutorials, and consuming screen reader feedback and other audio applications. On the other hand, BVI people expressed their desires to manipulate sounds and proposed ad-hoc solutions, such as distributing sound sources to different devices to allow the consumption of multiple sounds in parallel, adjusting sound characteristics to make sounds distinctive, or customizing existing sound libraries of applications. We synthesized our findings and proposed Sound Unblending to increase users’ perception and awareness of sounds in mixed reality, which includes six sound manipulation techniques:
  1. (1)
    Ambience Builder, which controls the ambient sounds or noises, to shift the sense of presence between RW and VR by manipulating acoustic transparency to occlusive opacity.
  2. (2)
    Feature Shifter, which modifies sound characteristics, such as the volume, pitch, and duration.
  3. (3)
    Spatializer, which controls the locations of sound sources to enhance the sense of immersion.
  4. (4)
    Stylizer, which manipulates a sound’s fidelity and overall style through different sound filters (e.g., high, low pass, human, robotic, anime).
  5. (5)
    Earcon Generator, which appends earcons for corresponding sound events, and
  6. (6)
    Prioritizer, which prioritizes multiple sounds by manipulating time (e.g., delaying less important sounds), location (e.g., left and right), and sound characteristics (e.g., the more important, the higher the volume).
We prototyped these sound manipulations in three Unity-simulated scenarios, focusing on tasks along the Reality-Virtuality Continuum (Milgram and Kishino, 1994). First, in a RW-Focused scenario when navigating a busy street, we explore ways to integrate virtual sounds when users primarily focus on RW tasks. Second, in a VR-Focused scenario, we explore ways to integrate real-world sounds when users primarily focus on VR tasks when listening to an audiobook. Third, in a Fully-Mixed scenario, we explore ways to interweave RW and VR sounds to enhance users’ perception of both realities when attending a hybrid conference.
@@
To understand how Sound Unblending manipulations can affect the perception of MR sounds, we conducted a user study (Odom et al., 2012a, 2014) with eighteen BVI participants who experienced the three simulated scenarios and identified sounds. In each scenario, we applied pre-defined sound manipulations on real-world and virtual sounds to enhance the perception of the mixed-reality soundscape. We compared Sound Unblending with the other two ends of the auditory Reality-Virtuality continuum (Milgram and Kishino, 1994; Larsson et al., 2010; McGill et al., 2020): full acoustic transparency that enhances the presence of the real world, and noise cancellation that enhances the immersion of the virtual reality (Figure 1a).
@@
Our results showed that sound manipulations significantly improved BVI people’s ability to perceive and manage sound information compared to full transparency and noise cancellation in our three simulated scenarios. The conditions also had varying effects on participants’ performance across the three scenarios. Sound Unblending also significantly reduced participants’ cognitive load in perceiving and managing sounds compared to full transparency and noise cancellation. Additionally, participants shared ideas for how they would further customize the sound manipulations in each scenario. Our evaluations demonstrate that sound unblending effectively improved MR sound awareness for BVI people.
@@
To demonstrate the generalizability and practicality of Sound Unblending, we further developed three real-world example applications based on our content analysis on online posts and user customization mentioned in our study: (i) an audio-adaptive online meeting web application that addresses the conflict between screen reader sounds and meeting conversations, (ii) a mixed-reality content-aware image exploration application that provides stylized and spatialized audio feedback based on real-world and virtual content, and (iii) a mobile navigation application that analyzes real-world and virtual sound events to identify opportune moments for delivering audio directions.
@@
  • The concept of Sound Unblending to make MR sound awareness accessible for BVI people, through six sound manipulators derived from our content analysis on BVI forums.
@@
  • Results from a user study with eighteen BVI people to provide empirical evidence of significantly enhanced sound awareness through sound unblending manipulations.
  • Three real-world example applications to demonstrate the generalizability and practicality of Sound Unblending.
@@ 2. Related work
Our work was motivated and situated from prior works on mixed reality, MR accessibility, and soundscape personalization.
@@ 2.2. Soundscape Formation and Personalization
Augmented Audio Reality (AAR) integrates virtual sounds into RW soundscapes. Larsson et al. (Larsson et al., 2010) posited that AR or AV in the concept of the Reality-Virtuality continuum can also adapt to auditory domains. McGill et al. (McGill et al., 2020) found that acoustically transparent headphones increase the sense of presence of reality while noise-canceling headphones diminish it. Advancements in noise-canceling technologies (also known as soundscaping technologies (Hagood, 2011)), have transformed the headphone user experience and emerged as a significant research field (Johansen et al., 2022). However, Haas et al. (Haas et al., 2018) described the limitations of soundscaping technologies: “current personal audio technology is not designed in a way that it allows users to handle their social context satisfactorily. Furthermore, information acquisition is made more difficult and users often have to make a choice between the surrounding acoustic environment and their own content and media.” Though technologies did not fully support soundscape curation, users still expressed desires to steer their soundscape (Haas et al., 2018) (e.g., blocking certain unwanted sounds, masking environmental sounds) and researchers also prototyped systems to adjust and adapt sounds based on users’ contexts (Cliffe et al., 2021; Mamuji et al., 2005). The absence of visual information creates the reliance on sounds for BVI people, which also reveals different needs and potential factors for curating soundscapes in different contexts (Rychtarikova, 2015).
Rather than altering the entire soundscape (e.g., acoustic transparency and occlusion), as facilitated by contemporary headphone technologies, in our work, we delve into more fine-grained sound manipulations. We seek to understand how these manipulations could enhance sound awareness for BVI people in various MR scenarios, and how BVI people would perceive and customize these sound manipulations.
@@ 3.1. Method
We conducted a content analysis from active online forums within the BVI community. First, two researchers communicated synchronously over Zoom and reviewed posts and their comments in online forums, including AppleVis (app, 2023b), and the Blind and Visually Impaired Community on Reddit (bli, 2023a). We started by filtering the posts by keywords such as “sound”, “audio”, “headphone”, “challenge”, and “scenario”, with a time range from 2021 to 2023. Then, we reviewed over 1000 posts and eliminated off-topic posts, such as social activity recruitment, debugging devices, and casual conversations. Ultimately, we collected 100 posts and comments highly associated with the difficulties or needs of sound consumption. Then, we used thematic analysis (Clarke et al., 2015) to analyze the data. Together, we reviewed and coded the 100 posts and comments in an online spreadsheet and had discussions to reach an agreement on the major themes described below.
@@ 3.2.2. Retaining Real-World Awareness in Noisy Environments
As mentioned, consuming both RW and VR sounds is common and inevitably creates several challenges for BVI people. RW noisy environments can overwhelm BVI people due to the difficulty in discerning important sounds out of noises, as one consumed Siri’s sound effects: “[…] they are difficult to hear in noisy environments even if you don’t have a hearing problem.” Although noise-cancellation headphones can block out noises, BVI people still want to retain awareness of their surroundings: “I want to be able to hear clearly what is around me at the same time I’m listening through them.” Bone-conduction headphones represent a potential alternative; however, they risk having VR sounds overshadowed by RW counterparts. Instead, one suggested dynamically adjusting virtual sound volume in accordance with the prevailing levels of RW sounds: “You would always be aware of ambient sound while you were wearing them, and I liked the idea that volume could be set to adjust automatically depending on the noise around you.” This idea echoes prior research on developing audio-adaptive systems for contextualized interactions (Cliffe et al., 2021; Mamuji et al., 2005; Zimmermann and Lorenz, 2008).
@@ 3.3. Summary
From our findings, we found different scenarios where BVI people encountered complex audio information. BVI people proposed their desires and solutions to better consume the complex audio information, including manipulating ambient noises while retaining real-world awareness, adjusting sound characteristics to make important sounds distinctive, distributing sound sources in different devices or locations, and having customizations on earcons or voice feedback in devices. These findings inspired the Sound Unblending manipulations described in the next section.
4. Sound Unblending Manipulations for Accessible MR Awareness
@@
Sound Unblending is a concept to enhance mixed-reality sound awareness, which includes six sound manipulations: Ambience Builder, Feature Shifter, Earcon Generator, Prioritizer, Spatializer, and Stylizer. We hypothesize that Sound Unblending manipulations can enhance the awareness of sounds for BVI people in mixed-reality environments. In this section, we describe how content analysis results and prior works inspired each sound manipulator.
@@
Ambience Builder modulates the ambient sounds or noises to shift presence between RW and VR by varying acoustic transparency and occlusion. Real-world noises often interfere with important RW or VR sounds (section 3.2.1 and 3.2.2). Also, environmental noises can enhance real-world grounding and presence in acoustic transparency as so-called “primitive hearing” (Ramsdell, 1978) that increases one’s presence in acoustic transparency than occlusion (McGill et al., 2020). This creates opportunities to balance the presence and awareness between RW and VR (Gilkey and Weisenberger, 1995; Murray et al., 2000; Larsson et al., 2010). The concept is similar to unreleased Apple Adaptive Audio (app, 2023a) by interpolating different auditory transparency and noise cancellation (Figure 1d.1).
@@
Feature Shifter modifies sound characteristics. Fundamental characteristics of sounds, such as pitch, volume, and duration, impact sensation and perception in daily listening (Gaver and Norman, 1988; Ballou, 1987). Modifying these characteristics, as discussed in section 3.2.3, helps make sounds distinguishable. As shown in Figure 1d.2, Feature Shifter increases the volume of specific sound sources over ambient noises.
@@
Spatializer controls the locations of sound sources to enhance the sense of immersion. Spatial audio has been a long-standing research field in VR and an essential component to embody the sense of presence and immersion (Lombard and Ditton, 1997; Brinkman et al., 2015; André et al., 2012). In sections 3.2.4, BVI people also distributed sound sources for selective attention, facilitating awareness through different directions. As shown in Figure 1d.5, Spatializer places sound sources at different spatial locations.
@@
Stylizer manipulates a sound’s fidelity and overall style through different sound filters. Section 3.2.5 highlights BVI individuals’ suggestions for screen reader voice customization beyond the standard synthesis. This includes diverse sound filters (e.g., high, low pass, human, robotic, anime) to cater to varied preferences. As shown in Figure 1d.3, Stylizer uses a low pass filter to soften sharp drilling noises.
Earcon Generator appends earcons for corresponding sound events. Earcons (Blattner et al., 1989) signal object updates and deliver key audio information with minimal user attention. In complex MR environments, earcons can signal RW and VR events to reduce cognitive load instead of describing everything using speech. As shown in Figure 1d.3, an earcon indicates potential danger.
@@
Prioritizer manages sound prioritization by adjusting time, location, and characteristics. Overlapping sounds are common (sections 3.2.1, 3.2.2, 3.2.3). Unlike Feature Shifter and Spatializer, Prioritizer dynamically contrasts sound sources, making them distinct through spatial or characteristic manipulation. It also controls sound timing to prevent overlap, such as pausing or delaying virtual audio in MR when it conflicts with RW sounds. As shown in Figure 1d.4, Prioritizer delays audio directions during drilling noises.
@@
Prior research indicated that full acoustic transparency (FT) increases RW presence and awareness, while noise cancellation (NC) enhances VR presence and awareness (McGill et al., 2020) (Figure 1a). We, therefore, posit them as the two ends of auditory Reality and Virtuality Continuum (Larsson et al., 2010) for optimally augmenting sound awareness in their respective realms. Consequently, our study compared Sound Unblending manipulations (SU) with these two established conditions, hypothesizing that participants can achieve the best performance with SU in mixed-reality settings by combining the best of both worlds. Specifically, we aim to understand the following research questions in this study:
@@
    How are participants’ experiences and ways to further customize their soundscape for each scenario?

### v2

# SoundShift: Exploring Sound Manipulations for Accessible Mixed-Reality Awareness

@@ Abstract.
Mixed-reality (MR) soundscapes blend real-world sound with virtual audio from hearing devices, presenting intricate auditory information that is hard to discern and differentiate. This is particularly challenging for blind or visually impaired individuals, who rely on sounds and descriptions in their everyday lives. To understand how complex audio information is consumed, we analyzed online forum posts within the blind community, identifying prevailing challenges, needs, and desired solutions. We synthesized the results and propose SoundShift for increasing MR sound awareness, which includes six sound manipulations: Transparency Shift, Envelope Shift, Position Shift, Style Shift, Time Shift, and Sound Append. To evaluate the effectiveness of SoundShift, we conducted a user study with 18 blind participants across three simulated MR scenarios, where participants identified specific sounds within intricate soundscapes. We found that SoundShift increased MR sound awareness and minimized cognitive load. Finally, we developed three real-world example applications to demonstrate the practicality of SoundShift.
@@
journalyear: 2024copyright: acmlicensedconference: Designing Interactive Systems Conference; July 1–5, 2024; IT University of Copenhagen, Denmarkbooktitle: Designing Interactive Systems Conference (DIS ’24), July 1–5, 2024, IT University of Copenhagen, Denmarkdoi: 10.1145/3643834.3661556isbn: 979-8-4007-0583-0/24/07ccs: Human-centered computing Interaction design theory, concepts and paradigmsccs: Human-centered computing Mixed / augmented realityccs: Human-centered computing Virtual reality
@@
Figure 1. We present SoundShift, a concept to manipulate sounds to improve mixed-reality awareness. (a) SoundShift is situated in the auditory Reality-Virtuality Continuum with full transparency and noise cancellation as two ends, and comprises (b) six sound manipulators, which are Transparency Shift, Envelope Shift, Position Shift, Style Shift, Sound Append, and Time Shift. (c) In a scenario, a BVI user navigates a busy street with a white cane and audio directions. They sometimes may receive ringtones and pass by construction sites with drilling noises. (d.1) Transparency Shift makes the auditory transparency half to suppress nuanced noises while retaining real-world awareness. (d.2) Envelope Shift increases the white cane sounds to make them distinctive. (d.3) Sound Append plays an earcon to signal the danger, and Style Shift applies a low-pass filter to make drilling noise less sharp to hear. (d.4) Time Shift delays the audio directions when they conflict with drilling noises. (d.5) Position Shift places ringtone on the right and audio directions on the left to increase distinguishability.
[…]
Figure 1 An overview of SoundShift system Figure 1a In the lower part of the figure, SoundShift is in the middle, and Full Transparency and Noise Cancellation are at the two ends on the left and right sides. In the upper part of the figure, there is text visual domain as the main topic, and virtuality is pointed to the right while reality is pointed to the left. Also, there are some related works below it, such as RealityCheck, One Reality, Remixed Reality, A Dose of Reality and ModularHMD. Figure 1b SoundShift include six sound manipulators: Transparency Shift, Envelope Shift, Style Shift, Position Shift, Time Shift, and Sound Append. Figure 1c A screenshot of Unity 3D world. A man wearing a black jacket is navigating the street with the white cane. On the street, there are cars whizzing by and crowds talking to each other. Figure 1d This is the figure showing sound waveforms of three study conditions: SoundShift, Full Transparency, and Noise Cancellation. Details can be referred to the image captions.
@@ 1. Introduction
To understand the current practices in consuming complex audio information, we first conducted a content analysis from active online forums within the BVI community, where we collected posts and comments about the prevailing scenarios, challenges, and potential solutions for consuming complex audio information. We found that many scenarios require the consumption of RW [^1] RW sounds refer to sounds from RW environments, such as crowd, speaker, television. and VR [^2] VR sounds refer to sounds rendered through users’ hearing devices. sounds, such as navigating a busy street with audio directions, playing an instrument following music tutorials, and consuming screen reader feedback and other audio applications. On the other hand, BVI people expressed their desires to manipulate sounds and proposed ad-hoc solutions, such as distributing sound sources to different devices to allow the consumption of multiple sounds in parallel, adjusting sound characteristics to make sounds distinctive, or customizing existing sound libraries of applications. We synthesized our findings and proposed SoundShift to enhance users’ perception and awareness of sounds in mixed reality environments. This approach incorporates six sound manipulation techniques: Transparency Shift, Envelope Shift, Position Shift, Style Shift, Time Shift, and Sound Append.
@@
To understand how SoundShift manipulations can affect the perception of MR sounds, we conducted a user study with eighteen BVI participants who experienced the three simulated scenarios and identified sounds, including a RW-Focused scenario when navigating a busy street, a VR-Focused scenario when focusing on audiobook, and a Fully-Mixed scenario when attending a hybrid conference. In each scenario, we applied pre-defined sound manipulations on real-world and virtual sounds to enhance the perception of the mixed-reality soundscape. We compared SoundShift with the other two ends of the auditory Reality-Virtuality continuum (Milgram and Kishino, 1994; Larsson et al., 2010; McGill et al., 2020): full acoustic transparency that enhances the presence of the real world, and noise cancellation that enhances the immersion of the virtual reality (Figure 1a).
@@
Our results showed that sound manipulations significantly improved BVI people’s ability to perceive and manage sound information compared to full transparency and noise cancellation in our three simulated scenarios. The conditions also had varying effects on participants’ performance across the three scenarios. SoundShift also significantly reduced participants’ cognitive load in perceiving and managing sounds compared to full transparency and noise cancellation. Additionally, participants shared ideas for how they would further customize the sound manipulations in each scenario. Our evaluations demonstrate that SoundShift effectively improved MR sound awareness for BVI people.
@@
To demonstrate the generalizability and practicality of SoundShift, we further developed three real-world example applications based on our content analysis on online posts and user customization mentioned in our study: (i) an audio-adaptive online meeting web application that addresses the conflict between screen reader sounds and meeting conversations, (ii) a mixed-reality content-aware image exploration application that provides stylized and spatialized audio feedback based on real-world and virtual content, and (iii) a mobile navigation application that analyzes real-world and virtual sound events to identify opportune moments for delivering audio directions.
@@
  • The concept of SoundShift to make MR sound awareness accessible for BVI people, through six sound manipulators derived from our content analysis on BVI forums.
@@
  • Results from a user study with eighteen BVI people to provide empirical evidence of significantly enhanced sound awareness through SoundShift manipulations.
  • Three real-world example applications to demonstrate the generalizability and practicality of SoundShift.
@@ 2. Related work
Our work was motivated and situated from prior work on mixed reality, MR accessibility, and soundscape personalization.
@@ 2.2. Soundscape Formation and Personalization
Augmented Audio Reality (AAR) integrates virtual sounds into RW soundscapes and is prevalent in our everyday lives (Krzyzaniak et al., 2019). Larsson et al. (Larsson et al., 2010) posited that AR or AV in the concept of the Reality-Virtuality continuum can also adapt to auditory domains. McGill et al. (McGill et al., 2020) found that acoustically transparent headphones increase the sense of presence of reality while noise-canceling headphones diminish it. Advancements in noise-canceling technologies (also known as soundscaping technologies (Hagood, 2011)), have transformed the headphone user experience and emerged as a significant research field (Johansen et al., 2022). However, Haas et al. (Haas et al., 2018) described the limitations of soundscaping technologies: “current personal audio technology is not designed in a way that it allows users to handle their social context satisfactorily. Furthermore, information acquisition is made more difficult and users often have to make a choice between the surrounding acoustic environment and their own content and media.” Though prior research (Cliffe et al., 2021; Mamuji et al., 2005; Sawhney and Schmandt, 2000; Bederson, 1995) or the current commercial headphones support adapting the audio based on users’ environment, users still expressed desires to steer and personalize their soundscape in a fine-grained manner (Haas et al., 2018) (e.g., blocking certain unwanted sounds, masking environmental sounds). This is particularly crucial for BVI individuals, who may have different vision levels and reliance on auditory cues, resulting in varying needs for curating soundscapes in different contexts (Rychtarikova, 2015). Recently, advancements in AI within the auditory domain have enhanced AAR and opened avenues for tailoring soundscaping technologies to users’ needs. For example, several works have facilitated intelligent sound extraction (Veluri et al., 2023b, a; Chatterjee et al., 2022), enabling users to selectively filter or emphasize specific sounds. Furthermore, sounds can be intelligently adapted to the user’s activities, offering seamless experiences. This includes altering music based on driving contexts to improve in-car music experiences (Kari et al., 2021), or integrating audio notifications with ongoing music to reduce disruption and annoyance (Alexander Wang, 2024).
Inspired by these needs and trends in personalizing soundscapes, we explore more fine-grained sound manipulations tailored specifically for BVI individuals, who, as experts in audio technologies, may have a unique reliance and strategies on manipulating sounds. We seek to understand how these fine-grained manipulations could enhance sound awareness for BVI people in various MR scenarios, and how BVI people would perceive and customize these sound manipulations.
@@ 3.1. Method
We conducted a content analysis from active online forums within the BVI community. First, two researchers communicated synchronously over Zoom and reviewed posts and their comments in online forums, including AppleVis (app, 2023b), and the Blind and Visually Impaired Community on Reddit (bli, 2023a). We started by filtering the posts by keywords such as “sound”, “audio”, “headphone”, “challenge”, and “scenario.” We reviewed posts from 2023 backward until we had collected over 1000 posts, ultimately reaching back to 2021. Then, we eliminated off-topic posts, such as social activity recruitment, debugging devices, and casual conversations. Ultimately, we collected 100 posts and comments highly associated with the difficulties or needs of sound consumption. Then, we used thematic analysis (Clarke et al., 2015) to analyze the data. Together, we reviewed and coded the posts and comments in an online spreadsheet and had discussions to reach an agreement on the major themes described below.
@@ 3.2.2. Retaining Real-World Awareness in Noisy Environments
As mentioned, consuming both RW and VR sounds is common and inevitably creates several challenges for BVI people. RW noisy environments can overwhelm BVI people due to the difficulty in discerning important sounds out of noises, as one consumed Siri’s sound effects: “[…] they are difficult to hear in noisy environments even if you don’t have a hearing problem.” Although noise-cancellation headphones can block out noises, BVI people still want to retain awareness of their surroundings: “I want to be able to hear clearly what is around me at the same time I’m listening through them.” From the online forums, bone-conduction headphones are frequently noted as a potential alternative by BVI people; however, there is a risk that VR sounds may be overshadowed by their RW counterparts. Instead, one suggested dynamically adjusting virtual sound volume in accordance with the prevailing levels of RW sounds: “You would always be aware of ambient sound while you were wearing them, and I liked the idea that volume could be set to adjust automatically depending on the noise around you.” This idea echoes prior research on developing audio-adaptive systems for contextualized interactions (Cliffe et al., 2021; Mamuji et al., 2005; Zimmermann and Lorenz, 2008).
@@ 3.3. Summary
Our findings revealed different scenarios where BVI people encountered complex audio information. These scenarios had different focuses spanning across the real world (e.g., traffic), virtual sounds (e.g., screen reader), or a combination of both (e.g., listening to music from the hearing device and playing instruments). Furthermore, BVI people proposed their desires and solutions to better consume the complex audio information, including manipulating ambient noises while retaining real-world awareness, adjusting sound characteristics to make important sounds distinctive, distributing sound sources in different devices or locations, and having customizations on earcons or voice feedback in devices. These findings inspired the three mixed-reality scenarios in our study and SoundShift manipulations described in the next section.
4. SoundShift Manipulations for Accessible Mixed Reality Awareness
@@
SoundShift is a concept to enhance mixed-reality sound awareness, which includes six sound manipulations: Transparency Shift, Envelope Shift, Sound Append, Time Shift, Position Shift, and Style Shift. We hypothesize that SoundShift manipulations can enhance the awareness of sounds for BVI people in mixed-reality environments. In this section, we describe how content analysis results and prior works inspired each sound manipulator.
@@
Transparency Shift modulates the ambient sounds or noises to shift presence between RW and VR by varying acoustic transparency. Acoustic transparency is a common description for headphones that blend virtual audio with real-world sounds (McGill et al., 2020). Real-world noises can enhance the sense of real-world grounding and presence, as so-called “primitive hearing” (Ramsdell, 1978). An increased sense of presence was also found when a user wears acoustically transparent headphones than the noise-cancellation ones (McGill et al., 2020). This creates opportunities to balance the presence and awareness between RW and VR through the degree of acoustic transparency (Gilkey and Weisenberger, 1995; Murray et al., 2000; Larsson et al., 2010). The concept is similar to Apple Adaptive Audio (app, 2023a) by interpolating different auditory transparency and noise cancellation (Figure 1d.1).
@@
Envelope Shift modifies the dynamic aspects of sounds, such as volume or pitch, affecting their perceived loudness and tonal characteristics over time. Fundamental characteristics of sounds, such as pitch, volume, and duration, impact sensation and perception in daily listening (Gaver and Norman, 1988; Ballou, 1987). Modifying these characteristics, as discussed in section 3.2.3, helps make sounds distinguishable. As shown in Figure 1d.2, Envelope Shift increases the volume of specific sound sources over ambient noises.
@@
Position Shift controls the locations of sound sources to enhance the sense of immersion. Spatial audio has been a long-standing research field in VR and an essential component to embody the sense of presence and immersion (Lombard and Ditton, 1997; Brinkman et al., 2015; André et al., 2012). In sections 3.2.4, BVI people also distributed sound sources for selective attention, facilitating awareness through different directions. As shown in Figure 1d.5, Position Shift places sound sources at different spatial locations.
@@
Style Shift transforms the timbre and fidelity of sounds through various filters to modify their aesthetic and emotional impact. Section 3.2.5 highlights BVI individuals’ suggestions for screen reader voice customization beyond the standard synthesis. This includes diverse sound filters (e.g., high, low pass, human, robotic, anime) to cater to varied preferences. As shown in Figure 1d.3, Style Shift uses a low pass filter to soften sharp drilling noises.
@@
Time Shift adjusts the timing of sounds to prioritize or deprioritize them within a soundscape, controlling auditory focus. Overlapping sounds are common (sections 3.2.1, 3.2.2, 3.2.3). Time Shift controls sound timing to prevent overlap, such as pausing or delaying virtual audio in MR when it conflicts with RW sounds. As shown in Figure 1d.4, Time Shift delays audio directions during drilling noises.
Sound Append appends earcons for corresponding sound events. Earcons (Blattner et al., 1989) signal object updates and deliver key audio information with minimal user attention. In complex MR environments, earcons can signal RW and VR events to reduce cognitive load instead of describing everything using speech. As shown in Figure 1d.3, an earcon indicates potential danger.
@@
Prior research indicated that full acoustic transparency (FT) increases RW presence and awareness, while noise cancellation (NC) enhances VR presence and awareness (McGill et al., 2020) (Figure 1a). We, therefore, posit them as the two ends of auditory Reality and Virtuality Continuum (Larsson et al., 2010) for optimally augmenting sound awareness in their respective realms. Consequently, our study compared SoundShift manipulations with these two established conditions, hypothesizing that participants can achieve the best performance with SS in mixed-reality settings by combining the best of both worlds. Specifically, we aim to understand the following research questions in this study:
@@
    How do participants describe their experiences and ways to further customize their soundscape for each scenario?